# OnePerc: A Randomness-aware Compiler for Photonic Quantum Computing

### v2

@@ 6.1. FlexLattice IR
Specifically, a FlexLattice graph is defined by the following features.
  1. (1)
    The graph has a (2+1)-D structure, with each 2D layer being a subgraph of a lattice with a fixed size.
  2. (2)
    Edges on the third dimension connect nodes at the same positions across different 2D lattice layers, irrespective of whether these layers are adjacent or not.
  3. (3)
    Each node has at most two edges in the third dimension, i.e., at most one with nodes on preceding layers and at most one with those on subsequent layers.
@@ 6.2. Instruction Set
An IR graph state can be implemented on the virtual hardware by a low-level instruction set defined as below, with nodes in the high-level program graph state represented by g_node and nodes on the virtual hardware represented by v_node. If not adjusted by the low-level instructions, qubits in the physical graph state are subject to $Z$-measurements by default. A virtual node can be mapped to a g_node or used as an acilla node for routing. In the former case the corresponding qubit will be measured according to the measurement pattern, while in the latter case it will be measured in $X$- or $Y$-basis adaptively. A virtual node can also be stored into or retrieved from a virtual cache by pushing or poping its surrounding qubits to or from the delay lines. Spatial edges between adjacent nodes on the same layer can be enabled by setting associated qubits to $X$- or $Y$-measurements, while temporal edges between nodes at the same position of adjacent layers can be enabled by establishing an adjacent-layer time-like connection.
@@
Cross-layer time-like connections between layer $m$ and layer $n\enspace(>m)$ can be established by a combination of three instructions, i.e., storing the node at layer $m$ into the virtual cache, retrieve it at layer $n-1$, and enable a temporal edge between layer $n-1$ and layer $n$. For example, the cross-layer temporal edge between ancilla node $A$ at (1,1,0) and graph node $A_{1}$ (1,1,2) in Fig. 9(d) can be implemented with the instructions below. Note that retrieving v_node at layer $n-1$ does not conflict with the original v_node at layer $n-1$. This is because the original node at layer $n-1$ would not have an edge with layer $n$, since each node in a FlexLattice graph has at most one edge with preceding layers. This implies that the original node will either have no further edges or will be stored in the virtual cache at layer $n-1$.
map_v_node((1, 1, 0), A)
@@
make_v_node_ancilla((1, 1, 2))
enable_temporal_v_edge((1, 1, 1), (1, 1, 2)
6.3. Offline Mapping
With this IR, graph state mapping algorithms such as that in OneQ can be utilized as an offline pass to produce an efficient IR program. To support large-scale programs, we extend OneQ’s mapping algorithm with two further optimizations. First, to reserve enough space for routing and avoid node congestion, we limit the percentage of incomplete nodes on each layer within a certain threshold, with incomplete nodes defined as those mapped nodes whose edges are not all mapped yet. Second, to mitigate the increasing demand on classical memory for graph information storage, we employ a refresh mechanism, which periodically retrieves all nodes stored in the virtual cache, refreshing them by connecting them with current one or multiple layers, and storing them again.
@@ 7.1. Experiment Setup
We compare the performance of our framework with the efficient photonic MBQC compiler OneQ. Since OneQ is not able to handle fusion failures, we employ it with a repeat-until-success strategy. Specifically, for each RSL we conduct the fusions implied by OneQ repeatedly until all fusions are successful. Subsequently, the successful RSL is fused with its preceding RSLs. If failures occur in the inter-RSL fusions, the entire compilation is reset and repeated until success.
@@
Aligning with OneQ, we evaluate the performance of compilation with two metrics: the number of consumed RSL, denoted by #RSL, and the number of required fusions, denoted by #fusion In particular, a smaller #RSL indicates less execution time of the program and less chance for photon loss, while a smaller #fusion implies less operations and less chance for error occurrence.
Table 1. Benchmark Programs.
Fusion Success Rate
#Qubits
Virtual Hardware Size
RSL Size
0.90
4
2x2
24x24
9
3x3
36x36
25
5x5
60x60
0.75
4
2x2
48x48
25
5x5
120x120
64
8x8
192x192
@@
We adopt the same photonic hardware architecture with OneQ, as introduced in Section 2. In the main experiment, the comparison with OneQ is performed with 4-qubit star-like resource states, and the sizes of hardware for different benchmarks are listed in Table 1. Subsequent experiments for further analysis are conducted with 7-qubit star-like resource states, which naturally have sufficient degrees for forming 3D lattice-like graph states.
@@
We select a set of benckmark programs including Quantum Approximate Optimization Algorithm (QAOA), Quantum Fourier transform (QFT), Ripple-Carry Adder (RCA) (RCA, ) and Variational Quantum Eigensolver (VQE). For QAOA, we choose the graph maxcut problem on randomly generated graphs. Specifically, the graphs are generated by randomly connecting half of all its possible edges. For VQE, we follow the commonly used full-entanglement ansatz, which proves to be an expressive ansatz (qiskit_vqe, ; vqe_ansatz, ). In table 1, we list the benchmarks with their numbers of qubits in the circuit representation. We also list the sizes of virtual hardware layers, along with the required sizes of RSLs to construct them.
@@
Fusion Success Rate
Benchmark Name
OneQ #RSL
OnePerc #RSL
#RSL Improv.
OneQ #Fusion
OnePerc #Fusion
#Fusion Improv.
0.90 (hyper-advanced)
QAOA-4
304
84
3.62
13,990
117,664
0.12
QFT-4
3,759
174
21.59
180,634
274,155
0.66
RCA-4
3,107
237
13.11
63,814
373,646
0.17
VQE-4
56
22
2.55
1,707
33,526
0.05
QAOA-9
$>10^{6}$
240
$>10^{3}$
$>10^{10}$
855,354
$>10^{4}$
QFT-9
570
$>10^{3}$
2,031,813
$>10^{3}$
RCA-9
1,017
$>10^{2}$
3,627,950
$>10^{3}$
VQE-9
156
$>10^{3}$
555,065
$>10^{4}$
QAOA-25
768
$>10^{3}$
7,637,711
$>10^{3}$
QFT-25
2,418
$>10^{2}$
24,065,102
$>10^{2}$
RCA-25
3,111
$>10^{2}$
30,962,172
$>10^{2}$
VQE-25
705
$>10^{3}$
7,010,656
$>10^{3}$
0.75 (practical)
QAOA-4
1,708
48
35.58
119,731
169,431
0.71
QFT-4
$>10^{6}$
210
$>10^{3}$
$>10^{10}$
746,977
$>10^{4}$
RCA-4
$>10^{6}$
201
$>10^{3}$
$>10^{10}$
714,835
$>10^{4}$
VQE-4
1,017
23
44.22
25,354
96,332
0.26
QAOA-25
$>10^{6}$
705
$>10^{3}$
$>10^{10}$
15,781,250
$>10$
QFT-25
2,271
$>10^{2}$
50,835,771
RCA-25
3,252
$>10^{2}$
72,795,212
VQE-25
759
$>10^{3}$
17,292,345
QAOA-64
2,787
$>10^{2}$
62,386,302
QFT-64
9,609
215,095,078
RCA-64
8,001
179,100,397
VQE-64
3,132
70,109,041
@@ 7.2. Experiment Result
In this subsection, we firstly show the performance of our compiler in comparison with OneQ. Then we analyze the effects of underlying resource states, hardware size and fusion success probability, for which we only focus on #RSL. This is because unlike OneQ, the #fusion in OnePerc is predictable from its #RSL, thus following a same trend with #RSL.
@@
Table 2 presents the comparison of our framework with OneQ. The results indicate a significant reduction of #RSL and #fusion by our framework. Specifically, the results show that OneQ can work only in the region of small programs and high fusion success probabilities. When the fusion success probability decreases to a practical value around 0.75, it takes more than $10^{6}$ RSLs even for executing the 4-qubit benckmarks. This implies OneQ’s non-scalability due to its lack of capability in systematically handling the randomness of fusion failure. In contrast, our framework can work well with a practical success probability, demonstrating an increasing outperformance over OneQ as the programs scale up.
@@
Resource State Size
Our compiler has a general applicability to the underlying resource states of various sizes. Fig. 10(a) illustrates the varying #RSL when executing the programs with star-like resource states of different sizes, i.e., consisting of different numbers of qubits. It can be seen that the #RSL decreases as the size of resource states increases. This is because a larger resource state can participate in fusions with more qubit degrees, without the need of increasing the degrees by merging multiple RSLs.
Hardware Size
Our compiler has an adaptability to various hardware sizes. Fig. 10(b) shows the varying #RSL when executing the programs on photonic hardware of different sizes, i.e., with different sizes of RSLs. It can be seen that a larger photonic hardware leads to a reduced #RSL, which indicates that our framework can effectively utilize the computing resource as it scales up. In particular, a larger RSL can enable a larger renormalized lattice, thus a larger virtual hardware. This provides the offline mapping with an increased space for flexible routing, thereby reducing the required logical layer and, consequently, the #RSL.
Fusion Success Probability
Our compiler has a capability of tolerating fusion failures at a practical level. Fig. 10(c) shows the varying #RSL when executing programs under different fusion success probabilities. It can be seen that our compiler can tolerate a fusion success probability as low as 0.66, with the #RSL decreasing with a higher fusion success probability. This is because a higher fusion success probability results in a larger renormalized lattice on RSLs, enabling a larger virtual hardware. This provides the offline mapping with an increased space for flexible routing, thereby reducing the required logical layer and, consequently, the #RSL.
7.3. Scalability and Parallelism
@@
Figure 11. Scalability and parallelism of OnePerc with 7-qubit resource states, with fusion success probability being 0.75 for (b)(c), average node size chosen as $20\times 20$ for (b)(c), hardware size being $100\times 100$ for (b) and $240\times 240$ for (c).
Scalability
Our compiler presents a great scalability, characterized by the stable overhead as the computing scales up. Fig. 11(a) shows the average node size of 2D renormalization as the hardware size increases. In particular, it keeps stable against the hardware size, being smaller with a higher fusion success probability. Fig. 11(b) shows the average ratio of RSL to logical layers as the program size increases. In particular, it first increases with the program size then soon gets stable at around 3, implying the successful formation of a logical layer about every 3 RSLs. These stabilities provide a predictability of the resource consumption and ensures the scalability of our framework.
Parallelism
While our framework achieves a notable real-time scalability with a modular 2D renormalization, it involves a tradeoff between the parallelism and the resource efficiency. This is because the presence of intervals between the modules reduces the available resource on each RSL (Fig. 8). To evaluate this resource overhead, Fig. 11(c) depicts the size of the renormalized 2D lattice against the number of modules. For comparison, the red dots represent the renormalized lattice size of a non-modular algorithm in an unlimited time, while the black dots represent the renormalization size of the non-modular algorithm in a time restricted by that consumed by the modular approach.
It can be seen that the renormalized lattice by the modular approach is around 60% of that by the non-modular approach with unlimited time, which decreases slightly with the number of modules. This is because an increased number of modules leads to a higher probability of being unable to connect the corresponding paths across different modules. However, the renormalized lattice is significantly larger than that can be achieved by the non-modular in the restricted time, which ranges from 2×2$\times$2 × to 7×7$\times$7 × as the number of modules increases from 4 to 16. This is very important since the time for the online algorithm is always restricted by the limited lifetime of photons. Overall, Fig. 11(c) indicates that our framework can significantly improve the real-time scalability with a relatively stable overhead of computing resource.

### v3

@@ 6.1. Virtual Hardware
6.2. FlexLattice IR and Offline Mapping
With this virtual hardware, graph state mapping algorithms such as that in OneQ can be utilized as an offline pass to enhance the efficiency of program execution. Specifically, the mapping onto virtual hardware transforms a program graph state to an equivalent IR program with compatible structure with the virtual hardware, which is referred to as a FlexLattice IR based on its structural features. This process is illustrated by Fig. 11(a) $\rightarrow$ (c) $\rightarrow$ (d).
To further improve the mapping efficiency and scale to larger programs, we extend OneQ’s mapping algorithm with three optimizations.
  • First, to map graph nodes as early as possible, we replace the static partition in OneQ with a dynamic scheduling. Specifically, we analyze the dependency among graph state qubits (measurement_calculus, ), representing it with a directed acyclic graph (DAG) and updating the front layer of the DAG as nodes are consumed by the mapping.
  • Second, to reserve enough space for routing and avoid node congestion, we enforce an upper-limit for the occupancy of incomplete nodes on each virtual hardware layer (25% by default), with incomplete nodes defined as those mapped nodes whose edges are not all mapped yet.
  • Third, to mitigate the increasing demand on classical memory for graph information storage, we propose a refresh mechanism, which periodically retrieves all nodes stored in the virtual memory, refreshing them by mapping onto multiple layers of the virtual hardware, and then storing them again.
@@ 6.3. Instruction Set
A FlexLattice IR program can be executed by transforming to a set of intermediate-level instructions, which guides the real-time physical operations to generate necessary connections among logical nodes through the reshaping algorithm in Section 5. By default, qubits in the physical graph state are subject to Z-measurements, which means that edges are disabled on the virtual hardware unless explicitly enabled by the intermediate-level instructions. We list the intermediate-level instructions as the following, with nodes in the high-level program graph state denoted as g_node and nodes on the virtual hardware denoted as v_node.
@@
By map_v_node() and make_v_node_ancilla(), a virtual node can be mapped by a g_node or used as an ancilla node to facilitate routing. In the former case, the physical qubit corresponding to v_node will be measured in the basis of the g_node, while in the latter case, it will be measured in $X$- or $Y$-basis to play as a wire (depending on whether the wire length is even or odd). By store_v_node() and retrieve_v_node(), a virtual node can also be stored into or retrieved from the virtual memory by pushing or poping its surrounding physical qubits to or from the delay lines. By enable_spatial_v_edge(), a spatial edge between adjacent nodes on the same layer can be enabled by setting associated qubits to $X$- or $Y$-measurements.
By enable_temporal_v_edge(), a temporal edge between logical nodes at the same coordinate of adjacent layers can be enabled. Establishment of a cross-layer edge between layer $m$ and layer $n\enspace(>m)$ can be realized through the combination of three instructions: storing the node at layer $m$ into the virtual memory, retrieving it at layer $n-1$, and enabling a temporal edge between layer $n-1$ and layer $n$. For example, the cross-layer temporal edge between ancilla node $A_{1}$ at (1,1,0) and graph node $A$ at (1,1,2) in Fig. 11(d) can be implemented with the instructions below. Note that retrieving v_node at layer $n-1$ does not conflict with the original v_node at layer $n-1$ (i.e., node $N$ in Fig. 11(d)). This is because the original node at layer $n-1$ would not have an edge with layer $n$, since each node in a FlexLattice IR has at most one edge with preceding layers. This implies that the original node will either have no further edges or will be stored in the virtual memory at layer $n-1$.
make_v_node_ancilla((1, 1, 0))
@@
enable_temporal_v_edge((1, 1, 1), (1, 1, 2))
map_v_node((1, 1, 2), A)
@@ 7.1. Experiment Setup
We compare the performance of our framework with the efficient photonic MBQC compiler OneQ. Since OneQ is not able to handle fusion failures, we employ it with a repeat-until-success strategy. Specifically, for each RSL we conduct the fusions instructed by OneQ repeatedly until all fusions are successful. Subsequently, the successful RSL is fused with its preceding RSLs. If failures occur in the inter-RSL fusions, the entire compilation is restarted and repeated until success.
Table 1. Benchmark Programs.
Fusion Success Rate #Qubits Virtual Hardware Size RSL Size
0.90 4 2x2 24x24
9 3x3 36x36
25 5x5 60x60
0.75 4 2x2 48x48
25 5x5 120x120
64 8x8 192x192
100 10x10 240x240
@@
Aligning with OneQ, we evaluate the performance of compilation with two metrics: the number of consumed RSL, denoted by #RSL, and the number of required fusions, denoted by #fusion. In particular, a smaller #RSL indicates less execution time of the program and less chance for photon loss, while a smaller #fusion implies less operations and less chance for error occurrence.
@@
We adopt the same photonic hardware architecture with OneQ, as introduced in Section 2. In the main experiment (Table 2, 3), the comparison with OneQ is performed on 4-qubit star-like resource states, with the sizes of hardware for different benchmarks listed in Table 1. Experiments for further analysis are conducted with 7-qubit star-like resource states, which naturally have sufficient degrees for forming 3D lattice-like graph states.
@@
We select a set of benckmark programs including Quantum Approximate Optimization Algorithm (QAOA), Quantum Fourier transform (QFT), Ripple-Carry Adder (RCA) (RCA, ) and Variational Quantum Eigensolver (VQE). For QAOA, we choose the graph maxcut problem on randomly generated graphs. Specifically, the graphs are generated by randomly connecting half of all its possible edges. For VQE, we follow the commonly used full-entanglement ansatz, which proves to be an expressive ansatz (qiskit_vqe, ; vqe_ansatz, ). In table 1, we list the benchmarks with their numbers of qubits in the circuit representation. We also list the sizes of virtual hardware layers, which are chosen to correspond with the qubit quantities, along with the required sizes of RSLs needed to generate them, which are determined through Fig. 16, as explained later.
@@
Fusion Success Rate Benchmark Name OneQ #RSL OnePerc #RSL #RSL Improv. OneQ #Fusion OnePerc #Fusion #Fusion Improv.
0.90 (hyper-advanced) QAOA-4 304 84 3.62 13,990 117,664 0.12
QFT-4 3,759 174 21.59 180,634 274,155 0.66
RCA-4 3,107 237 13.11 63,814 373,646 0.17
VQE-4 56 22 2.55 1,707 33,526 0.05
QAOA-9 $>10^{6}$ 240 $>10^{3}$ $>10^{10}$ 855,354 $>10^{4}$
QFT-9 570 $>10^{3}$ 2,031,813 $>10^{3}$
RCA-9 1,017 $>10^{2}$ 3,627,950 $>10^{3}$
VQE-9 156 $>10^{3}$ 555,065 $>10^{4}$
QAOA-25 768 $>10^{3}$ 7,637,711 $>10^{3}$
QFT-25 2,418 $>10^{2}$ 24,065,102 $>10^{2}$
RCA-25 3,111 $>10^{2}$ 30,962,172 $>10^{2}$
VQE-25 705 $>10^{3}$ 7,010,656 $>10^{3}$
0.75 (practical) QAOA-4 1,708 48 35.58 119,731 169,431 0.71
QFT-4 $>10^{6}$ 210 $>10^{3}$ $>10^{10}$ 746,977 $>10^{4}$
RCA-4 $>10^{6}$ 201 $>10^{3}$ $>10^{10}$ 714,835 $>10^{4}$
VQE-4 1,017 23 44.22 25,354 96,332 0.26
QAOA-25 $>10^{6}$ 882 $>10^{3}$ $>10^{10}$ 19,743,350 $>10$
QFT-25 2,271 $>10^{2}$ 50,835,771
RCA-25 3,252 $>10^{2}$ 72,795,212
VQE-25 759 $>10^{3}$ 17,292,345
QAOA-64 3,339 $>10^{2}$ 191,341,276
QFT-64 9,000 515,801,985
RCA-64 9,324 534,311,489
VQE-64 3,042 174,321,702
Table 3. Effect of refresh on the performance of OnePerc, considering 4-qubit resource state, a fusion success rate of 0.75, refresh rate of 50 logical layers, and 32GB of RAM.
Benchmark #Qubits Non-refreshed #RSL Refreshed #RSL
QAOA 25 882 999
64 - 4,284
100 - 8,325
QFT 25 2,271 2,637
64 - 9,945
100 - 19,494
RCA 25 3,252 3,870
64 - 10,206
100 - 16,056
VQE 25 759 774
64 - 3,555
100 - 7,551
@@ 7.2. Experiment Result
In this subsection, we first show the performance of our compiler in comparison with OneQ, then analyze the effects of underlying resource states, hardware size and fusion success probability, for which we only focus on the #RSL metric. This is because unlike OneQ, the #fusion in OnePerc is predictable from its #RSL, thus following a same trend with #RSL.
@@
Table 2 presents the comparison of our framework with OneQ. The results indicate a significant reduction of #RSL by our framework, as well as a significant reduction of #fusion when the circuits are beyond 4 qubits. Specifically, the experiments show that OneQ can work only in the region of small programs and high fusion success probabilities. When the fusion success probability decreases to a practical value around 0.75, it takes more than $10^{6}$ RSLs to even execute the 4-qubit benckmarks. This implies OneQ’s non-scalability due to its lack of capability in systematically handling the randomness of fusion failure. In contrast, our framework can work well with a practical success probability, demonstrating an increasing outperformance over OneQ as the programs scale up.
An obstacle of scaling up the experiments in Table 2 is the large classical memory required in the real-time stage for the storage of graph information. Indeed, the 64-qubit benchmarks in Table 2 takes a RAM as much as 192 GB. This can be overcome by the refresh mechanism proposed in Section 6, with an overhead of increased #RSL. Under the practical fusion success rate of 0.75, Table 3 shows the effect of refresh given 32 GB RAM. It can be seen that while the 32 GB RAM can only afford 25-qubit benchmarks without refresh, it allows for benchmarks of up to 100 qubits with a refresh every 50 logical layers. Compared with the performance of 25-qubit benchmarks (Table 2 or 3) and 64-qubit benchmarks (Table 2) without refresh, the introduction of refresh leads to an average increase of 15.6% in #RSL for 25-qubit benchmarks and an average increase of 13.3% in #RSL for 64-qubit benchmarks.
7.3. Sensitivity Analysis
Resource State Size
Our compiler has a general applicability to the underlying resource states of various sizes. Fig. 12(a) illustrates the varying #RSL when executing the programs with star-like resource states of different sizes, i.e., consisting of different numbers of photonic qubits. It can be seen that the #RSL decreases as the size of resource states increases. This is because a larger resource state can participate in fusions with more qubit degrees, without the need of increasing the degrees by merging multiple RSLs.
Hardware Size
Our compiler has an adaptability to various hardware sizes. Fig. 12(b) shows the varying #RSL when executing the programs on photonic hardware of different RSL sizes. It can be seen that a larger photonic hardware leads to a reduced #RSL, which indicates that our framework can effectively utilize the computing resource as it scales up. In particular, a larger RSL can enable a larger renormalized lattice, thus a larger virtual hardware. This provides the offline mapping with an increased space for flexible routing, thereby reducing the required logical layer and the #RSL.
Fusion Success Probability
Our compiler has a capability of tolerating fusion failures at a practical level. Fig. 12(c) shows the varying #RSL when executing the programs under different fusion success probabilities. It can be seen that our compiler can tolerate a fusion success probability as low as 0.66, with the #RSL decreasing as the fusion success probability increases. This is because a higher fusion success probability results in a larger renormalized lattice on RSLs, enabling a larger virtual hardware. This provides the offline mapping with an increased space for flexible routing, thereby reducing the required logical layer and the #RSL.
7.4. Scalability
Resource Consumption
Our compiler presents a great scalability in resource consumption, characterized by the stable overhead as the computing scales up. Fig. 13(a) shows the suitable average node size of 2D renormalization as the hardware size increases, corresponding to the average node size at which the renormalization success probability approaches 1 in Fig. 16. As can be seen, it keeps stable against the variation of hardware size, being smaller with a higher fusion success probability. Fig. 13(b) shows the average ratio of RSL to logical layers as the program size increases. It first increases with the program size and then soon gets stable at a value around 3, implying the successful formation of a logical layer about every 3 RSLs. These stable behaviours provide a predictability of the resource consumption and ensures the scalability of our framework.
@@ 7.4. Scalability
Modularity Overhead
The real-time scalability of our framework can be greatly enhanced through a modular 2D renormalization, which reduces the latency for each RSL by a factor corresponding to the modular number. However, this comes with an overhead, as the presence of intervals between the modules (as illustrated in Fig. 10) reduces the available resource on each RSL. To evaluate this resource overhead, Fig. 13(c) depicts the size of the renormalized 2D lattice against the number of modules, with the MI ratio (as defined in Fig. 10) ranging from 2 to 19. For comparison, the red dots represent the renormalized 2D lattice size by a non-modular algorithm in an unlimited time, while the black dots represent the renormalized size by the non-modular algorithm in a time restricted by that consumed by the modular approach.
@@
Figure 13. Scalability and parallelism of OnePerc with 7-qubit resource states. The node size $n\times n$ in (a) corresponds to the smallest node size where the renormalization success rate approaches to 1 in Fig.  16.
Figure 14. Oneline proccessing time for each RSL with 7-qubit resource states. RSL size is $96\times 96$ for (a); fusion success rate is 0.75 for (a)(b); average node size is chosen as $24\times 24$ for (a)(b); MI ratio is chosen as 7 for (b).
Figure 15. Offline compilation time on a virtual hardware, with the virtual hardware size being $4\times 4$ for (a). The virtual hardware sizes correspond to the sizes that can be formed by the RSL settings in Fig. 14.
It can be seen that the size of renormalized 2D lattice by the modular approach is around 60% of that by the non-modular approach with unlimited time (red), which decreases slightly with the number of modules. This is because an increased number of modules leads to a higher probability of being unable to connect the corresponding paths across different modules. However, the renormalized lattice is significantly larger than that can be achieved by the non-modular approach restricted in the same time (black), ranging from 2×2$\times$2 × to 6×6$\times$6 × as the number of modules increases from 4 to 16. This is very important since the time for the online algorithm is always restricted by the limited lifetime of photons. Overall, Fig. 13(c) indicates that the modular approach in our framework can significantly improve the real-time scalability with a reasonable overhead of computing resource.
Compilation Time
We show the online and offline compilation time of the benchmarks in Fig. 14 and Fig. 15, with the compiler implemented in Python. From Fig. 14(a) it can be seen that the online processing time for each RSL stays stable as the program size increases. From Fig. 14(b), which takes an average of all 36-qubit benchmarks, it can be seen that the processing time for each RSL increases with RSL size, but can be significantly reduced by employing a modular renormalization. For offline compilation time, Fig. 15(a) shows that it increases with the program size. Fig. 15(b) shows that it decreases with the virtual hardware size first and then increases. This occurs because an excessively small virtual hardware size leads to a significant total depth, whereas an overly large virtual hardware size results in extended compilation times for each logical layer.